OpenStudy (valiyuh):

How long will it take take light moving at 300,000 km/s to reach us from the sun? The sun is 15,000,000 km from earth.

OpenStudy (anonymous):

Just use dimensional analysis to see how to approach this. You have a speed in km/s. You have a distance in km. How could you use those two numbers to produce a time, with units of seconds? t = distance / speed = km / (km / s) = s

OpenStudy (anonymous):

\[Time = \frac{ Distance }{Speed}\] First of all..Here the distance between sun and earth is wrong..It is about: 149,600,000 FACT:THis distance is taken as 1 Astronomical unit; SPped of light 3 * 10 ^ 5 km Therefore Time is : \[\frac{ 1.49 * 10^8}{3 * 10 ^5}\] \[= \frac{ 1.49 * 10^3 }{ 3 }\] \[=\frac{ 149 * 10^3 }{3 * 10^2} = \frac{ 149 * 10^1 }{ 3} = 496 seconds= 496/60 = 8.2 minutes[Avg.]\]